# Intelligent Surfaces Empowered Wireless Network: Recent Advances and The Road to 6G

### 15 Jan 2024 v2

@@ II Reflection-based IS: Design Issues and State-of-the-art Results
TABLE II: A toy example of blind beamforming when $N=4$, $K=2$, and $T=6$.
@@ II-A Reflection Design
In order to fully exploit ISs’ potential, it requires judicious design of the reflection coefficients of all IS elements, leading to the passive reflection/beamforming optimization problem. Most of the existing design methods for IS passive reflection/beamforming are based on a two-stage procedure, i.e., the cascaded BS-IS-user channel is first estimated via downlink/uplink pilots and the reflection coefficients are then optimized accordingly [41, 39, 58, 59]. However, for current cellular protocols, the pilots are dedicated to the use of direct channel estimation between the BS and users, which, however, are insufficient to estimate the extra IS channels. Thus, the estimation of IS channels requires significant modifications of the current network protocol. Moreover, due to massive reflective elements (REs) of IS, a large number of additional pilots are needed to guarantee the successful recovery/estimation of high-dimensional channel matrices/vectors of the BS-IS, IS-user, or cascaded BS-IS-user channels, which entail excessively high overhead for existing wireless systems.
@@
To enable the application of ISs in current and future wireless communication systems, recent research efforts have been devoted to developing more practical approaches for IS passive reflection/beamforming design. Specifically, instead of relying on the baseband complex-valued pilots, the IS reflection coefficients can be optimized based on the received signal power measurement at the user terminals [60, 61, 62, 63, 64, 65], which are easy to acquire in existing cellular networks, e.g., from the reference signal received power (RSRP). Thus, the power measurement-based IS reflection/beamforming design does not require additional pilots and is fully compatible to the current network protocol. Under this paradigm, the new methods for IS reflection/beamforming design can be mainly classified into three categories, i.e., blind reflection/beamforming, channel recovery-based reflection/beamforming, and beam training, as shown in Fig. 5.
@@
Specifically, the blind reflection/beamforming approach optimizes the phase shift of each RE based on the user’s power measurement without the need of any explicit channel state information (CSI) estimation [60, 61, 66, 67]. In contrast, the channel recovery-based approach first estimates the cascaded BS-IS-user channel based on the user’s power measurement and then optimizes the IS reflection/beamforming accordingly [62, 63]. Note that the above two approaches are practically applicable to both sub-6 GHz frequency bands with rich multi-path channels and mmWave frequency bands with LoS-dominant channels. However, they are generally less efficient in mmWave frequency bands due to the channel/path sparsity and high path loss of mmWave signals. In particular, the received signal power may be small or even negligible if the IS reflection/beamforming vector is not aligned with the dominant channel path. Furthermore, the training overhead of these methods may be unnecessarily large because they do not fully exploit the inherent channel sparsity in mmWave frequency bands. To overcome the above issues, beam training serves as a more efficient approach for IS passive beamforming design in mmWave systems by selecting the best directional beam from a predefined beamforming codebook [64, 65]. In the following, the above three approaches for IS passive reflection/beamforming design are discussed in more details.
@@ II-A1 Blind Reflection/Beamforming
Let us start with a naive algorithm called Random-Max Sampling (RMS) for blind reflection/beamforming. Its idea is fairly simple: just try out a sequence of random samples of the phase shifts of IS and choose the one with the best performance thus far. The following toy example illustrates how RMS works. Assume that IS has four REs; denote by $\theta_{n}$ the phase shift of the $n$th RE, $n=1,2,3,4$; assume also that each phase shift $\theta_{n}$ takes on value from the binary set $\{0,\pi\}$. Suppose that a total of six random samples of the phase shift array $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ have been tested as shown in Table II. The performance of each random sample is evaluated by the utility value, e.g., we can take the received signal-to-noise ratio (SNR) (which is in direct proportion to the RSRP) as the utility for the single-antenna transmission. Observe that RMS gives the solution $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\pi,0,\pi,\pi)$ in this example because the fourth random sample has the highest utility value. Clearly, RMS can attain the global optimum so long as all the possible combinations of $\{0,\pi\}$ have been tested.
@@
However, a real-world IS has far more than four REs as assumed in the above toy example, e.g., there are 256 REs integrated into the IS prototype in [60], so it is already intractable to search through the solution space even if each phase shift has only two choices. Then a natural idea is to partially explore the solution space. But how much portion of the entire solution space should RMS explore in order to achieve “good” performance? To formalize the above question, we denote by $N$ the number of REs, $T$ the number of random samples, and $K$ the number of choices for each phase shift. For ease of discussion, we focus on the single-antenna transmission and use the SNR as the utility. A fundamental problem is how fast the SNR scales with $N$ if $T$ is much smaller than the solution space size $K^{N}$, which is termed as the SNR boost.
@@
The above fundamental problem has been answered in [60] for RMS. Roughly speaking, the authors in [60] showed that $\mathrm{SNR}\propto N\log T$ by the RMS method. Moreover, it has been shown in [68] that the scaling rate is at most quadratic in $N$, and this upper bound is tight in the ideal case where every reflected channel is perfectly aligned with the direct channel. Thus, in order to achieve a quadratic boost of SNR, we must set $T$ to be exponentially large in $N$ for the RMS method. However, $T$ can only be polynomially large in $N$ in practice. As a result, the actual performance of RMS is $\mathrm{SNR}\propto N\log N$. One may then have a follow-up question: Is it possible to strike a quadratic boost of SNR (which is the theoretical upper-bound as revealed in [68]) by using a polynomial number of random samples, $T$?
@@
Two recent works [60, 61] provide a positive answer to the above question by devising the so-called Conditional Sample Mean (CSM) method. In a nutshell, CSM first calculates the mean utility conditioned on each possible phase-shift choice for a particular RE, and then chooses the phase shift with the highest conditional mean utility for this RE. We still use the toy example in Table II to illustrate CSM. Let us see how $\theta_{1}$ is decided. Recall that there are two possible choices, 00 and $\pi$, for the phase shift; we first average out all the random samples with $\theta_{1}=0$ so as to obtain the mean utility conditioned on $\theta_{1}=0$, i.e., E[utilityθ_1=0] = 2.8+1.0+0.43=1.4, and similarly find the mean utility conditioned on $\theta_{1}=\pi$ as E[utilityθ_1=π] = 1.5+3.3+0.33=1.7. Comparing the above two conditional mean utility values, the CSM method chooses $\theta_{1}=\pi$. Likewise, the other phase shifts are determined as $\theta_{2}=0$, $\theta_{3}=\pi$, and $\theta_{4}=0$. Intuitively, CSM aims at the average goodness of setting $\theta_{1}$ to a particular possible choice while the rest phase shifts are still pending; with every phase shift optimized in this fashion independently, CSM enables an efficient configuration of IS.
@@ II-A2 Channel Recovery-Based Reflection/Beamforming
The blind reflection/beamforming design in general requires a sufficiently large number of random samples (i.e., IS phase-shift realizations) as well as corresponding received signal power measurements to achieve the upper bound on the SNR boost in IS-aided communication systems. This is fundamentally because the explicit CSI between the BS, IS, and user is not exploited, which results in the underutilization of power measurements as well as high training overhead. To overcome this limitation, a new channel recovery-based reflection/beamforming approach was proposed in [62, 63] based on the received signal power measurement as well. Specifically, by collecting the received signal power under different IS reflections, the cascaded BS-IS-user channel (or its autocorrelation matrix) is first estimated. Then, based on the obtained CSI, the IS reflection/beamforming coefficients are optimized more efficiently for improving the communication performance.
@@
As the phase information of the received signal is missing in its measured power, a fundamental question for the power measurement-based channel estimation is whether the complex-valued channel vector can be recovered by only exploiting the received signal powers which are positive real numbers. The results in [62] revealed that if the number of power measurements is sufficiently large, the autocorrelation matrix of the cascaded BS-IS-user channel can be uniquely determined. As such, the cascaded BS-IS-user channel can be recovered with an unknown phase ambiguity that is common to all IS REs. A special case occurs if the REs have only two phase-shift choices, i.e., 0 and $\pi$. In this case, the channel autocorrelation matrix cannot be distinguished with its conjugate. Nonetheless, since the passive beamforming vector of IS has real values only in this binary phase-shift case, such a phase ambiguity does not affect the SNR boost at the user’s receiver [62].
@@
In general, the channel recovery can be formulated as an optimization problem which aims to find a channel vector satisfying the power measurement constraints under different IS reflections. Since the power measurement constraints are non-convex, a useful way is to lift up the channel/beamforming vectors into matrices by calculating their autocorrelation matrices [62]. As a result, the power measurement constraints become linear to the channel autocorrelation matrix. Since the channel autocorrelation matrix has the rank of one, the optimal solution can be found by minimizing its rank, subject to the power measurement constraints. By applying proper relaxations, the rank-minimization problem can be solved by using the semidefinite programming (SDP)-based method [62]. To reduce the computational complexity as well as improve the robustness to noise, a neural network (NN)-based solution was proposed in [63] to estimate the cascaded BS-IS-user channel vector. Specifically, a single-layer NN was constructed by leveraging the structure whereby the channel and beamforming vectors impact the received signal power. The input and output of the NN are defined as the real and imaginary parts of the IS reflection vector and the expected received signal power, respectively. The weights of the single-layer NN correspond to the cascaded channel, which can be estimated by training the NN via minimizing the loss function, i.e., the mean square error (MSE) between the NN’s output and the measured signal power. For both SDP-based and NN-based channel recovery, the IS passive beamforming vector can be further optimized based on the estimated CSI. Various optimization techniques proposed in the literature can be utilized, such as semidefinite relaxation (SDR) [69], block coordinate descent (BCD) [70], successive convex approximation (SCA) [71], the penalty-based method [72], and the projected gradient ascent/descent method [73].
@@ II-A3 Beam Training
TABLE III: IS beam training design.
@@
Note that the above blind and channel recovery-based reflection/beamforming approaches can be applied in real time for improving the instantaneous channels or offline for improving the channel statistical distribution in sub-6 GHz bands. For mmWave and higher frequency bands, high beamforming gain is required to compensate for the severe path loss. Due to the channel/path sparsity, codebook-based beam training with directional beams is a more efficient approach to design IS beamforming without the need of explicit CSI estimation as illustrated in Fig. 8. In the following, we discuss efficient methods for IS beam training in the far- and near-field channel scenarios, respectively.
@@
3.1) Far-Field Beam Training: First, consider the case where both the BS and users are located in the far-field region of the IS as shown in Fig. 8(b). Instead of performing joint BS and IS beam training that incurs high training overhead, a more effective approach is conducting an offline BS beam training to establish a high-gain BS-IS link by exploiting the fixed locations of the BS and IS. Then, only the IS beam training needs to be performed in the online phase for selecting the best IS beam in a predesigned codebook that yields the strongest signal power at the user’s receiver. For narrow-band systems, sequential single-beam training is a straightforward method that exhaustively searches over all possible beam directions (i.e., codewords) at the IS. As revealed in [91, 74], there generally exists a fundamental tradeoff between maximizing the achievable throughput and minimizing the beam training overhead, i.e., a more accurate beam training leads to a higher passive beamforming gain, while at the cost of higher training overhead as well. To reduce the IS beam training overhead, one line of research has developed efficient IS beam codebooks to reduce the beam-search space. For example, the authors in [74] proposed a small-size IS codebook with wide beamwidth, which is achieved by a quadratic phase-shift design for enabling higher-order variations at IS phase shifts. In addition, under the IS discrete phase-shift constraint, an SCA-based algorithm was proposed in [75] to design the IS codebook with beamwidth-adjustable codewords, while at the expense of decreased beamforming gain. Alternatively, hierarchical IS codebook design and beam search is also an efficient way to reduce the beam training overhead without compromising much the beamforming gain [65, 76]. Specifically, a hierarchical IS beam training method was proposed in [65], which first identifies the best beam sector with wide IS beams and then refines the IS beam in the sector using narrow beams. However, the IS codewords may not be orthogonal in practice due to IS hardware imperfections, hence resulting in reduced beam training success rate. In addition, another hierarchical IS beam search method was proposed in [76], which uses two multi-mainlobe codewords in each layer for beam sweeping and exploits the beam discrimination in each layer to effectively increase the success rate. However, hierarchical beam training approaches generally require frequent user feedback to refine the beam selection, thus incurring high feedback/training overhead that scales with the number of users. To circumvent this issue, the authors in [64] proposed a new multi-beam training method, where IS REs are divided into multiple sub-arrays, which allows for simultaneous beam sweeping over different directions. Then, each user can determine its optimal IS beam direction by comparing the received signal power/SNR over time, thus greatly reducing the training overhead yet without the need for successive user feedback in IS-aided multi-user systems. Besides the pilot-based beam training, other approaches for IS beam training in narrow-band systems have also been recently proposed. For instance, the authors in [77] proposed a new binocular camera-mounted IS architecture, where the visual inear-fieldormation is collected to aid the IS in locating the BS/access point (AP) as well as dynamic beam tracking with the users, thus eliminating the need for dedicated training symbols and signal feedback. In [78], the authors considered an ISAC system with IS and proposed a simultaneous beam training and target sensing scheme. Specifically, the BS simultaneously senses the targets and performs the beam training for both the users and IS, based on which the LoS IS-user channel is determined by exploiting their geometry relationship.
@@
For wide-band systems, IS beam training designs become more challenging due to the lack of frequency selectivity in IS phase shifts; thus they remain the same over all frequencies. This results in the so-called beam squint effect, where the beams at different frequencies are generally dispersed at different directions. To address this issue, the authors in [79] proposed to first obtain the optimal IS phase shifts for each subcarrier and then determine the common phase shift by maximizing the upper bound of the sum achievable rate over all subcarriers in orthogonal frequency-division multiplexing (OFDM) systems. Alternatively, the authors in [80] proposed to equip each IS sub-surface with a time-delay unit to enable frequency-dependent phase shifts over the entire frequency band, thus offering a flexible tradeoff between wide-band communication performance and hardware design complexity.
@@
3.2) Near-Field Beam Training: Next, we consider the case where an extremely large-scale IS (XL-IS) is equipped with a huge number of REs to compensate for the product-distance path-loss. In this case, the BS and users are more likely to be located in the near-field region of the XL-IS as shown in Fig. 8(c). We still assume that the BS-IS link has been aligned offline thanks to the fixed BS/IS locations, and thus focus on the IS beam training with users only. Note that directly applying the existing IS beam training methods for far-field scenarios will inevitably incur degraded beam training performance due to the channel mismatch. Thus, it is necessary to develop new near-field IS beam training methods accounting for (radiative) near-field spherical wave-fronts [92, 93].
@@
To this end, several new codebooks catering for near-field beam training have been proposed, which can be applied for designing the near-field IS beam training. For example, a Cartesian-domain codebook was devised in [81], which covers the entire three-dimensional (3D) space by performing uniform sampling in the Cartesian coordinate system. However, the dimension of the Cartesian-domain codebook is determined by the product of the numbers of sampled points at the $x$/$y$/$z$ axis, which is practically very large and thus incurs excessive beam training overhead. To address this issue, a more efficient polar-domain codebook design was proposed in [82], where the near-field region is sampled with non-uniform angle-range grids by minimizing adjacent codeword correlation. Based on this polar-domain codebook, one straightforward near-field IS beam training method is by exhaustively searching over all codewords, which, however, requires a large number of training symbols proportional to the product of the numbers of sampled angles and ranges. To tackle this challenge, one efficient approach is by applying the two-phase near-field beam training method proposed in [83] for IS. Specifically, the XL-IS/array first estimates the user angle by performing conventional DFT-codebook based beam sweeping and exploiting the “angle-in-the-middle” effect in the received beam pattern. Then, given a set of estimated user angles, the XL-IS adopts the polar-domain beam training in the range domain to determine the user range. Numerical results showed that this two-phase method can significantly reduce the required near-field beam training overhead for XL-array/IS with negligible rate loss. This work was further extended in [84] that directly utilizes the conventional DFT-codebook to resolve the user angle and range by more effectively exploiting the received signal powers. Alternatively, the authors in [85] proposed to first determine the user range via a designed omnidirectional IS codeword, and then perform the near-field beam sweeping in the angle domain to obtain user angle information.

### 24 Mar 2024 v3

@@ II Reflection-based IS: Design Issues and State-of-the-art Results
TABLE III: A toy example of blind beamforming when $N=4$, $K=2$, and $T=6$.
@@ II-A Reflection Design
In order to fully exploit ISs’ potential, it requires judicious design of the reflection coefficients of all IS elements, thus giving rise to passive reflection/beamforming optimization problem [115, 116, 117, 118, 119, 120]. Most of the existing design methods for IS passive reflection/beamforming are based on a two-stage procedure, i.e., the cascaded BS-IS-user channel is first estimated via downlink/uplink pilots [121, 122, 123] and the reflection coefficients are then optimized accordingly [95, 92, 124, 125, 126]. However, under the current cellular protocols, the pilots are dedicated to the use of direct channel estimation between the BS and users without IS reflection. Thus, the estimation of IS channels requires significant modifications of the current cellular protocols. Moreover, due to massive reflective elements (REs) of IS, a large number of additional pilots are needed to estimate the high-dimensional BS-IS, IS-user, or cascaded BS-IS-user channels [121, 122, 123, 127], which entail excessively high overhead for the existing wireless systems.
@@
To enable the seamless application of ISs in current and future wireless communication systems, recent research efforts have been devoted to developing more practical approaches for IS passive reflection/beamforming design. Specifically, instead of relying on the baseband complex-valued pilots, the IS reflection coefficients can be optimized based on the received signal power measurement at the user terminals [128, 59, 129, 130, 131, 132], which are easily accessible in the existing cellular networks, e.g., from the reference signal received power (RSRP). Thus, the power measurement-based IS reflection/beamforming design does not require additional pilots and is fully compatible to the current cellular protocols. Currently, such designs can be mainly classified into three categories, i.e., blind reflection/beamforming, channel recovery-based reflection/beamforming, and beam training, as shown in Fig. 5.
@@
Specifically, the blind reflection/beamforming approach optimizes the phase shift of each RE based on the user’s power measurement without the need of any explicit channel state information (CSI) estimation [128, 59, 133, 134]. In contrast, the channel recovery-based approach first estimates the cascaded BS-IS-user channel based on the user’s power measurement and then optimizes the IS reflection/beamforming accordingly [129, 130]. Note that the above two approaches are practically applicable to both sub-6 GHz bands with rich multi-path channels and mmWave frequency bands with LoS-dominant channels. However, they are generally less efficient in mmWave frequency bands due to the channel/path sparsity and severe path loss [^3] For each IS element, the channel condition (e.g., channel sparsity and number of multi-paths) is generally determined by the signal propagation environment and carrier frequency only, which is not affected by the aperture size of the whole IS. Nonetheless, the end-to-end effective channel of the BS-IS-user link is highly dependent on the IS aperture and beamforming/reflection design for both the instantaneous channels and statistical channel distribution.. In particular, the received signal power may not be sufficient for CSI recovery if the IS reflection/beamforming vector is not aligned with the dominant channel path. Furthermore, the training overhead of these methods may be unnecessarily large because they do not fully exploit the inherent channel sparsity in mmWave frequency bands. To overcome the above issues, beam training serves as a more efficient approach for IS passive beamforming design in mmWave systems by selecting the best directional beam from a predefined beamforming codebook [131, 132]. In the following, the above three power-measurement-based approaches for IS passive reflection/beamforming design are discussed in more details.
@@ II-A1 Blind Reflection/Beamforming
The blind IS reflection/beamforming approach applies to the discrete phase shift of REs. Let us start with a naive algorithm called Random-Max Sampling (RMS) for blind reflection/beamforming. Its idea is fairly simple: just try out a sequence of random samples of the phase shifts of IS and choose the one with the best performance thus far. The following toy example illustrates how RMS works. Assume that IS has four REs; denote by $\theta_{n}$ the phase shift of the $n$th RE, $n=1,2,3,4$; assume also that each phase shift $\theta_{n}$ takes on value from the binary set $\{0,\pi\}$. Suppose that a total of six random samples of the phase shift array $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})$ have been tested as shown in Table III. The performance of each random sample is evaluated by the utility value, e.g., we can take the received signal-to-noise ratio (SNR) (which is in direct proportion to the RSRP) as the utility for the single-antenna transmission. Observe that RMS gives the solution $(\theta_{1},\theta_{2},\theta_{3},\theta_{4})=(\pi,0,\pi,\pi)$ in this example because the fourth random sample has the highest utility value. Clearly, RMS can attain the global optimum so long as all the possible combinations of $\{0,\pi\}$ have been tested.
@@
However, a real-world IS has far more than four REs as assumed in the above toy example, e.g., there are 256 REs integrated into the IS prototype in [128], so it is already intractable to search through the solution space even if each phase shift has only two choices. Then a natural idea is to partially explore the solution space. But how much portion of the entire solution space should RMS explore in order to achieve “good” performance? To formalize the above question, we denote by $N$ the number of REs, $T$ the number of random samples, and $K$ the number of choices for each phase shift. For ease of discussion, we focus on the single-antenna transmission and use the SNR as the utility. A fundamental problem is how fast the SNR scales with $N$ if $T$ is much smaller than the solution space size $K^{N}$, which is termed as the SNR boost [^4] Note that the SNR boost refers to the scaling order of the SNR with respect to $N$, which is identical for both discrete and continuous phases shifts of IS..
@@
The above fundamental problem has been answered in [128] for RMS. Roughly speaking, the authors in [128] showed that the SNR boost behaves like $\mathrm{SNR}\propto N\log T$ by the RMS method. Moreover, it has been shown in [115] that the scaling rate is at most quadratic in $N$, and this upper bound is tight in the ideal case where every reflected channel is perfectly aligned with the direct channel. Thus, in order to achieve a quadratic boost of SNR, we must set $T$ to be exponentially large in $N$ for the RMS method. However, $T$ can only be polynomially large in $N$ in practice. As a result, the actual performance of RMS is $\mathrm{SNR}\propto N\log N$. One may then have a follow-up question: Is it possible to strike a quadratic boost of SNR (which is the theoretical upper-bound as revealed in [115]) by using a polynomial number of random samples, $T$?
@@
Two recent works [128, 59] provide a positive answer to the above question by devising the so-called Conditional Sample Mean (CSM) method; a more recent work [135] further extends CSM to the joint configuration of multiple ISs. In a nutshell, CSM first calculates the mean utility conditioned on each possible phase-shift choice for a particular RE, and then chooses the phase shift with the highest conditional mean utility for this RE. We still use the toy example in Table III to illustrate CSM. Let us see how $\theta_{1}$ is decided. Recall that there are two possible choices, 00 and $\pi$, for the phase shift; we first average out all the random samples with $\theta_{1}=0$ so as to obtain the mean utility conditioned on $\theta_{1}=0$, i.e., E[utilityθ_1=0] = 2.8+1.0+0.43=1.4, and similarly find the mean utility conditioned on $\theta_{1}=\pi$ as E[utilityθ_1=π] = 1.5+3.3+0.33=1.7. Comparing the above two conditional mean utility values, the CSM method chooses $\theta_{1}=\pi$. Likewise, the other phase shifts are determined as $\theta_{2}=0$, $\theta_{3}=\pi$, and $\theta_{4}=0$. Intuitively, CSM aims at the average goodness of setting $\theta_{1}$ to a particular possible choice while the rest phase shifts are still pending; with every phase shift optimized in this fashion independently, CSM enables an efficient configuration of IS.
@@ II-A2 Channel Recovery-Based Reflection/Beamforming
The blind reflection/beamforming design in general requires a large number of random samples (i.e., IS phase-shift realizations) as well as corresponding received signal power measurements to achieve the upper bound on the SNR boost in IS-aided communication systems. This is fundamentally because the explicit CSI between the BS, IS, and user is not exploited, which results in the underutilization of power measurements as well as high training overhead. To overcome this limitation, a new channel recovery-based reflection/beamforming approach was proposed in [129, 130] based on the received signal power measurement as well. Specifically, by collecting the received signal power under different IS reflections, the cascaded BS-IS-user channel (or its autocorrelation matrix) is first estimated. Then, based on the obtained CSI, the IS reflection/beamforming coefficients are optimized more efficiently for improving the communication performance. Note that the channel recovery-based approach applies to both discrete and continuous phase shifts of REs.
@@
As the phase information of the received signal is missing in its measured power, a fundamental question for the power measurement-based channel estimation is whether the complex-valued channel vector can be recovered by only exploiting the received signal powers which are positive real numbers. The results in [129] revealed that if the number of power measurements is sufficiently large, the autocorrelation matrix of the cascaded BS-IS-user channel can be uniquely determined. As such, the cascaded BS-IS-user channel can be recovered with an unknown phase ambiguity that is common to all IS REs. A special case occurs if the REs have only two phase-shift choices, i.e., 0 and $\pi$. In this case, the channel autocorrelation matrix cannot be distinguished with its conjugate. Fortunately, since the passive beamforming vector of IS has real values only in this binary phase-shift case, such a phase ambiguity does not affect the SNR boost at the user’s receiver [129].
@@
In general, the channel recovery can be formulated as an optimization problem which aims to find a channel vector satisfying the power measurement constraints under different IS reflections. Since the power measurement constraints are non-convex, one viable approach is by lifting up the channel/beamforming vectors into matrices by calculating their autocorrelation matrices [129]. As a result, the power measurement constraints become linear in the channel autocorrelation matrix. Since the channel autocorrelation matrix has the rank of one, the optimal solution can be found by minimizing its rank, subject to the power measurement constraints. By applying proper relaxations, the rank-minimization problem can be solved by using the semidefinite programming (SDP)-based method [129]. On the other hand, to reduce the computational complexity as well as improve the robustness to noise, a neural network (NN)-based approach was proposed in [130] to estimate the cascaded BS-IS-user channel vector. Specifically, it was revealed that the received signal power can be viewed as a single-layer NN, with its input and output corresponding to the real and imaginary parts of the IS reflection vector and the expected received signal power, respectively. The weights of the single-layer NN correspond to the cascaded channel, which can be estimated by training the NN via minimizing the loss function, i.e., the mean square error (MSE) between the NN’s output and the measured signal power. For both SDP-based and NN-based channel recovery, the IS passive beamforming vector can be further optimized based on the estimated CSI. Various optimization techniques proposed in the literature can be utilized, such as semidefinite relaxation (SDR) [136], block coordinate descent (BCD) [137], successive convex approximation (SCA) [138], the penalty-based method [139], and the projected gradient ascent/descent method [140].
@@ II-A3 Beam Training
TABLE IV: IS beam training design.
@@
Note that the above blind and channel recovery-based reflection/beamforming approaches can be applied in real time for improving the instantaneous channels or offline for improving the channel statistical distribution in sub-6 GHz bands. For mmWave and higher frequency bands, high beamforming gain is required to compensate for the severe path loss. Due to the channel/path sparsity, codebook-based beam training with directional beams is a more efficient approach to design IS beamforming without the need of explicit CSI estimation as illustrated in Fig. 8. In the following, we discuss efficient methods for IS beam training in the far- and near-field channel scenarios, respectively, under the assumption of continuous RE phase shifts; while these methods can be readily extended to the case of discrete RE phase shifts.
@@
3.1) Far-Field Beam Training: First, consider the case where both the BS and users are located in the far-field region of the IS as shown in Fig. 8(b). Instead of performing joint BS and IS beam training that incurs high training overhead, a more effective approach is conducting an offline BS beam training to establish a high-gain BS-IS link by exploiting the fixed locations of the BS and IS. Then, only the IS beam training needs to be performed in the real-time phase for selecting the best IS beam in a predesigned codebook that yields the strongest signal power at the user’s receiver. For narrow-band systems, sequential single-beam training is a straightforward method that exhaustively searches over all possible beam directions (i.e., codewords) at the IS. As revealed in [158, 141], there generally exists a fundamental tradeoff between maximizing the achievable throughput and minimizing the beam training overhead, i.e., a more accurate beam training leads to a higher passive beamforming gain, while at the cost of higher training overhead as well. To reduce the IS beam training overhead, one line of research has developed efficient IS beam codebooks to reduce the beam-search space. For example, the authors in [141] proposed a small-size IS codebook with wide beamwidth, which is achieved by a quadratic phase-shift design for enabling higher-order variations at IS phase shifts. In addition, under the IS discrete phase-shift constraint, an SCA-based algorithm was proposed in [142] to design the IS codebook with beamwidth-adjustable codewords, while at the expense of decreased beamforming gain. Alternatively, hierarchical IS codebook design and beam search is also an efficient way to reduce the beam training overhead without compromising much the beamforming gain [132, 143]. Specifically, a hierarchical IS beam training method was proposed in [132], which first identifies the best beam sector with wide IS beams and then refines the IS beam in the sector using narrow beams. However, the IS codewords may not be orthogonal in practice due to IS hardware imperfections, hence resulting in a reduced beam training success rate. In addition, another hierarchical IS beam search method was proposed in [143], which uses two multi-mainlobe codewords in each layer for beam sweeping and exploits the beam discrimination in each layer to effectively increase the success rate. However, hierarchical beam training approaches generally require frequent user feedback to refine the beam selection, thus incurring high feedback/training overhead that scales with the number of users. To circumvent this issue, the authors in [131] proposed a new multi-beam training method, where IS REs are divided into multiple sub-arrays, which allows for simultaneous beam sweeping over different directions. Then, each user can determine its optimal IS beam direction by comparing the received signal power/SNR over time, thus greatly reducing the training overhead yet without the need for successive user feedback in IS-aided multi-user systems. Besides pilot-based beam training, other approaches for IS beam training in narrow-band systems have also been recently proposed. For instance, the authors in [144] proposed a new binocular camera-mounted IS architecture, where the visual information is collected to aid the IS in locating the BS/access point (AP) as well as dynamic beam tracking with the users, thus eliminating the need for dedicated training symbols and signal feedback. In [145], the authors considered an ISAC system with IS and proposed a simultaneous beam training and target sensing scheme. Specifically, the BS simultaneously senses the targets and performs the beam training for both the users and IS, based on which the LoS IS-user channel is determined by exploiting their geometry relationship.
@@
For wide-band systems, IS beam training designs become more challenging due to the lack of frequency selectivity in IS phase shifts; thus they remain the same over all frequencies. This results in the so-called beam squint effect, where the beams at different frequencies are generally dispersed in different directions. To address this issue, the authors in [146] proposed to first obtain the optimal IS phase shifts for each subcarrier and then determine the common phase shift by maximizing the upper bound of the sum achievable rate over all subcarriers in orthogonal frequency-division multiplexing (OFDM) systems. Alternatively, the authors in [147] proposed to equip each IS sub-surface with a time-delay unit to enable frequency-dependent phase shifts over the entire frequency band, thus offering a flexible tradeoff between wide-band communication performance and hardware design complexity.
@@
3.2) Near-Field Beam Training: Next, we consider the case where an extremely large-scale IS (XL-IS) is equipped with a huge number of REs to compensate for the product-distance path-loss. In this case, the BS and users are more likely to be located in the near-field region of the XL-IS as shown in Fig. 8(c). We still assume that the BS-IS link has been aligned offline thanks to the fixed BS/IS locations, and thus focus on the IS beam training with users only. Note that directly applying the existing IS beam training methods for far-field scenarios will inevitably incur degraded beam training performance due to the channel mismatch. Thus, it is necessary to develop new near-field IS beam training methods accounting for (radiative) near-field spherical wave-fronts [159, 160, 161].
@@
To this end, several new codebooks catering to near-field beam training have been proposed, which can be applied to designing the near-field IS beam training. For example, a Cartesian-domain codebook was devised in [148], which covers the entire three-dimensional (3D) space by performing uniform sampling in the Cartesian coordinate system. However, the dimension of the Cartesian-domain codebook is determined by the product of the numbers of sampled points at the $x$/$y$/$z$ axis, which is practically very large and thus incurs excessive beam training overhead. To address this issue, a more efficient polar-domain codebook design was proposed in [149], where the near-field region is sampled with non-uniform angle-range grids by minimizing adjacent codeword correlation. Based on this polar-domain codebook, one straightforward near-field IS beam training method is by exhaustively searching over all codewords, which, however, requires a large number of training symbols proportional to the product of the numbers of sampled angles and ranges. To tackle this challenge, one efficient approach is by applying the two-phase near-field beam training method proposed in [150] for IS. Specifically, the XL-IS/array first estimates the user angle by performing conventional DFT-codebook based beam sweeping and exploiting the “angle-in-the-middle” effect in the received beam pattern. Then, given a set of estimated user angles, the XL-IS adopts the polar-domain beam training in the range domain to determine the user range. Numerical results showed that this two-phase method can significantly reduce the required near-field beam training overhead for XL-array/IS with negligible rate loss. This work was further extended in [151] that directly utilizes the conventional DFT-codebook to resolve the user angle and range by more effectively exploiting the received signal powers. Alternatively, the authors in [152] proposed to first determine the user range via a designed omnidirectional IS codeword, and then perform the near-field beam sweeping in the angle domain to obtain user angle information.